### v2

# Neural Stereo Video Compression with Hybrid Disparity Compensation

Abstract
Disparity compensation represents the primary strategy in stereo video compression (SVC) for exploiting cross-view redundancy. These mechanisms can be broadly categorized into two types: one that employs explicit horizontal shifting, and another that utilizes an implicit cross-attention mechanism to reduce cross-view disparity redundancy. In this work, we propose a hybrid disparity compensation (HDC) strategy that leverages explicit pixel displacement as a robust prior feature to simplify optimization and perform implicit cross-attention mechanisms for subsequent warping operations, thereby capturing a broader range of disparity information. Specifically, HDC first computes a similarity map by fusing the horizontally shifted cross-view features to capture pixel displacement information. This similarity map is then normalized into an “explicit pixel-wise attention score” to perform the cross-attention mechanism, implicitly aligning features from one view to another. Building upon HDC, we introduce a novel end-to-end optimized neural stereo video compression framework, which integrates HDC-based modules into key coding operations, including cross-view feature extraction and reconstruction (HDC-FER) and cross-view entropy modeling (HDC-EM). Extensive experiments on SVC benchmarks, including KITTI 2012, KITTI 2015, and Nagoya, which cover both autonomous driving and general scenes, demonstrate that our framework outperforms both neural and traditional SVC methodologies.
I Introduction
As stereoscopic cameras become increasingly common in newly launched smart products, such as virtual reality headsets and autonomous vehicles, the demand for efficient storage and transmission of large volumes of stereo video is growing. Stereo video compression (SVC) [1, 2, 3, 4] has recently gained significant attention for its ability to encode stereo videos into compact bitstreams, reducing both storage and transmission costs. However, developing an effective SVC algorithm remains a non-trivial research problem. A key challenge is reducing cross-view redundancy, where the disparity compensation is central to this effort. By aligning stereo views, disparity compensation enables the extraction of missing details and the reuse of shared information, which enhances reconstruction quality and improves compression efficiency. Traditional SVC [2, 1] rely on handcrafted disparity compensation methods, but these approaches are limited by the constraints of manually designed structures. The recent success of deep neural networks (DNNs) in stereo visual data processing [5, 6, 7, 8, 9, 10, 11] has paved the way for neural-based disparity compensation mechanisms. These modern techniques offer a more flexible and accurate solution for reducing cross-view redundancy in SVC.
Existing neural-based disparity compensation methods can be classified into two categories. The first category leverages explicit disparity information to align cross-view features. This information is typically derived from standard operations such as varying disparity shifts [12, 13, 3], predicted disparity shifts [14, 3], or a homography matrix [15]. Incorporating these explicit operations provides a clear structural prior with direct geometric guidance, thereby simplifying optimization. However, while this approach reduces the complexity of the search space by predominantly focusing on local pixel neighborhoods, its reliance on local feature matching can limit robustness in occluded regions and may lead to erroneous matches. On the other hand, another category produces implicit disparity information by applying a cross-attention mechanism across both views to generate an implicit similarity map [16, 17, 18]. Leveraging global matching, this approach dynamically aggregates cross-view features based on attention maps, thereby extending its range and better addressing corner cases such as noises and occlusions. However, the global matching operation required to produce the similarity map is typically harder to optimize and computationally more expensive than directly generating pixel displacement, resulting in slower convergence, increased training overhead, and greater optimization difficulty.
In this work, we introduce a hybrid disparity compensation (HDC) strategy that leverages the simplicity of explicit pixel displacement as a robust prior for straightforward optimization while harnessing the high performance of implicit cross-attention mechanisms for effective disparity compensation. Specifically, HDC begins by applying horizontal shifts to the features from both perspectives, generating two feature volumes corresponding to different disparity shifts. An element-wise dot product between these feature volumes computes a similarity map, which is normalized to produce an “attention score”. This score is then used in a cross-attention mechanism for each view, implicitly aligning the projected information from one view to the other to complete the disparity compensation process. Compared to conventional explicit compensation algorithms, our approach facilitates a more flexible search for disparity compensation by leveraging the dynamic nature of the attention score to encompass broader ranges. Additionally, our method streamlines the optimization process relative to implicit compensation techniques by employing standard operations (i.e., shifting) to extract cross-view disparity information, thereby providing a structural prior that enhances neural network learning.
Building upon HDC, we propose a novel end-to-end optimized neural stereo video compression (NSVC) framework that integrates HDC-based modules into key coding operations: cross-view feature extraction and reconstruction (HDC-FER) and cross-view entropy modeling (HDC-EM). Specifically, HDC-FER extracts and reconstructs enhanced cross-view features by aligning features from one view to another, thereby integrating information from both views to facilitate the encoding and decoding of each specific view. Additionally, HDC-EM improves the contexture coding procedure by partitioning the quantized latent features from each view into sub-quantized latent features along the channel dimension and then progressively and iteratively entropy encoding them. In this process, previously encoded sub-quantized latent features from both views are used as priors, and the HDC mechanism aligns sub-features from another view to the current coding feature, thereby refining these priors.
We conduct extensive experiments to evaluate our framework on widely used automotive stereo video benchmarks, including KITTI 2012 [19] and KITTI 2015 [20]. In contrast to previous NSVC methods [4, 3]—which have been evaluated solely on automotive scenes—we also test our approach on general multi-view scenes using the Nagoya dataset [21], a standard stereo video benchmark in MV-HEVC [1, 2]. Experimental results highlight the effectiveness of our framework: it surpasses both neural and traditional SVC methods in automotive scenarios and achieves performance comparable to MV-HEVC in general multi-view scenes. Our contributions can be summarized as follows:
  • We propose a hybrid disparity compensation, i.e., HDC, strategy that leverages explicit pixel displacement as a robust disparity prior and incorporates implicit cross-attention for improved compensation performance.
  • We develop an end-to-end optimized NSVC framework by integrating our HDC into key components, including cross-view feature extraction and reconstruction and cross-view entropy modeling.
  • Extensive experiments demonstrate that our framework outperforms existing SVC methods on autonomous driving benchmarks and achieves performance comparable to the standard MV-HEVC in general multi-view scenes.
II Related Work
II-A Video compression
Traditional video codecs [22, 23, 24] employ hand-crafted motion compensation modules to exploit temporal dependencies across frames. Despite their proven effectiveness, these hand-crafted modules face inherent limitations in optimizing the rate-distortion trade-off.
Recently, NVC has seen rapid advancements. Mainstream frameworks can be broadly categorized into residual coding [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] and conditional coding paradigms [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. Specifically, residual coding methods explicitly subtract predicted frames from the current frame. For instance, Lu et al. [51] introduced the first end-to-end NVC framework by incorporating learnable motion estimation and compensation. These schemes depend strongly on precise spatial alignment, as it is critical for effective explicit motion compensation. On the other hand, another category conditional coding, captures temporal dependencies directly within the latent feature space, enabling more efficient and adaptable compression. In particular, the DCVC family [37, 39, 41, 43, 45, 50] adopts advanced conditional coding architectures and has demonstrated significant improvements in video compression performance across a range of benchmarks.
While these methods achieve impressive results in single-view video compression, they are not readily applicable to MVC task due to their lack of consideration for cross-view redundancy inherent in multi-view settings, thereby limiting potential performance gains. To address this limitation, we propose a novel disparity compensation module and build upon DCVC-TCM [50], a state-of-the-art yet relatively lightweight member of the DCVC family, to strike a balance between compression performance and efficiency.
II-B Stereo Image Compression
Stereo image compression (SIC) techniques typically build upon conventional image compression methods by additionally leveraging cross-view redundancy through disparity compensation. This compensation process aligns stereo image pairs to improve both reconstruction quality and coding efficiency. Traditional SIC algorithms implement the disparity mechanism using hand-crafted modules such as non-separable vector lifting [52], wavelet [53], and variable-block [54]. However, similar to traditional video compression methods, such hand-crafted modules are inherently limited in their ability to support end-to-end optimization.
Recently, neural SIC approaches have achieved tremendous success by employing deep learning-based disparity compensation strategies, which can be broadly categorized into two types: explicit and implicit alignment. Specifically, explicit alignment methods [12, 14, 15, 55] utilize predefined disparity shifts to warp features across views, leveraging strong geometric priors such as DSIC [12], HESIC [15], DispSIC [14], and MASIC [55]. For example, Deng et al. [15] applies homography-based warping to align stereo features based on estimated planar transformations. Nevertheless, these methods rely heavily on accurate local correspondences, making them vulnerable to occlusions and disparity estimation errors.
On the other hand, implicit alignment methods [16, 17, 18, 13, 16] perform global feature interaction using cross-attention mechanisms to generate implicit similarity maps. For example, SASIC [13] and ECSIC [17] adopt attention-based matching strategy that enhances robustness and effectively handles occlusions and noise. However, generating such implicit similarity maps is computationally intensive during both training and inference, leading to slower convergence and increased training costs.
Although both categories have achieved notable success in the SIC domain, extending those SIC strategies to SVC presents a non-trivial challenge. In SVC, both performance and efficiency are critical, as disparity compensation must be applied to every stereo image pair at each time step. Poor performance or high computational cost can lead to significant error accumulation and increased latency. To address this, we propose a hybrid compensation strategy that leverages the strengths of both paradigms by combining the simplicity and robustness of explicit alignment with the high representational capacity of implicit cross-attention mechanisms. This hybrid design enables efficient optimization while maintaining high disparity compensation accuracy across time steps.
Figure 1: Left: Overview of our Neural Stereo Video Compression (NSVC) framework. At time step $t$, the input stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ is compressed and reconstructed into $\{\bm{\hat{x}}_{t}^{L},\bm{\hat{x}}_{t}^{R}\}$ using five key coding components. This process is conditioned on the previously reconstructed frame pair $\{\bm{\hat{x}}_{t-1}^{L},\bm{\hat{x}}_{t-1}^{R}\}$ and the corresponding feature pair $\{\bm{F}_{t-1}^{L},\bm{F}_{t-1}^{R}\}$ from time step t1t-1 as in established neural video compression (NVC) methods [50]. Middle: In the two key components, Motion Compression and Context Compression, we apply and extend our Hybrid Disparity Compensation (HDC) strategy. Specifically, HDC-FER mechanisms are used within their encoders and decoders to enhance the intermediate feature pair $\{\bm{K}_{t}^{L},\bm{K}_{t}^{R}\}$, while HDC-EM mechanisms are used within their entropy model to improve the quantized latent feature pair $\{\bm{\hat{y}}_{t}^{L},\bm{\hat{y}}_{t}^{R}\}$. Right: High-level syntax of the proposed HDC strategy. We first apply explicit pixel displacement to construct a similarity map, which is then used in a cross-attention mechanism to implicitly align features across views.
II-C Multi-view Video Compression
MVC techniques aim to compress video sequences captured from multiple synchronized viewpoints. Unlike single-view video coding, MVC simultaneously leverages temporal redundancy within each view and inter-view redundancy across different views through joint motion and disparity compensation. For instance, established standards such as MVC [1] and MV-HEVC [2] extend the conventional H.264 [22] and H.265 [23] architectures by incorporating inter-view reference frames and hierarchical prediction structures.
SVC serves as a foundational technique for those MVC frameworks. To enhance SVC performance, two NSVC approaches have emerged, leveraging learning-based disparity compensation techniques. Specifically, LSVC [4] employs a sequential compression scheme with deformable-based disparity warping, while LLSS [3] introduces a novel Bishift operation to construct cost volumes for parallel disparity modeling. To achieve high efficiency, both methods adopt explicit disparity shift-based mechanisms. However, as discussed in Sec. II-B, although explicit disparity compensation is computationally efficient, its performance is often constrained by limited robustness to errors and occlusions. In this work, we propose a hybrid compensation strategy that balances efficiency and accuracy, enabling high-performance SVC.
III Methodology
III-A Overall Architecture
Given a stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ at time step $t$, the objective of NSVC is to represent this pair using fewer bits via an encoding process, such that the original data can be accurately reconstructed through a corresponding decoding process, while preserving a specified level of visual quality. In this work, we present an enhanced stereo video codec built upon a modified version of the DCVC-TCM video compression framework [50], which is adapted to efficiently compress stereo video frames. An overview of the proposed framework is illustrated in the left part of Fig. 1.
Motion Estimation. Given the input stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ at time step $t$ and the reconstructed stereo frame pair $\{\bm{\hat{x}}_{t-1}^{L},\bm{\hat{x}}_{t-1}^{R}\}$ from the previous time step (retrieved from the decoded buffer), the motion estimation network initially estimates the motion vector pair $\{\bm{mv}_{t}^{L},\bm{mv}_{t}^{R}\}$.
Motion Compression. The estimated motion vector pair $\{\bm{mv}_{t}^{L},\bm{mv}_{t}^{R}\}$ is subsequently compressed by the motion compression network and reconstructed as $\{\bm{\hat{mv}}_{t}^{L},\bm{\hat{mv}}_{t}^{R}\}$. The motion compression network consists of an encoder, an entropy model, and a decoder, which are responsible for feature extraction, entropy modeling, and feature reconstruction, respectively. The encoder transforms the input motion vectors into a latent feature pair, which is then quantized and losslessly compressed using arithmetic coding. The entropy model estimates the probability distribution of the quantized latent feature pair to facilitate cross-entropy coding. The decoder reconstructs the motion vectors from the quantized latent features. During the feature extraction (resp., reconstruction) process, an intermediate feature pair is produced after each downsampling (resp., upsampling) operation.
Motion Compensation. The motion compensation network generates multi-scale context features for each view using the feature pair $\{\bm{F}_{t-1}^{L},\bm{F}_{t-1}^{R}\}$ from the previous time step, retrieved from the decoded buffer. It then utilizes the motion vector pair $\{\bm{\hat{mv}}_{t}^{L},\bm{\hat{mv}}_{t}^{R}\}$ to align the multi-scale context features to the current time step, producing an aligned multi-scale context feature pair $\{\bm{Ctx}^{L}_{1\sim 3,t},\bm{Ctx}^{R}_{1\sim 3,t}\}$.
Context Compression. The stereo frame pair $\{\bm{x}_{t}^{L},\bm{x}_{t}^{R}\}$ is compressed by the context compression network and reconstructed as $\{\bm{F}_{t}^{L},\bm{F}_{t}^{R}\}$, with the multi-scale context feature pair $\{\bm{Ctx}^{L}_{1\sim 3,t},\bm{Ctx}^{R}_{1\sim 3,t}\}$ serving as conditional input. This context feature pair is generated by a multi-scale temporal context mining network and incorporated into the feature extraction (resp., reconstruction) process to effectively exploit temporal information, following the design in [50]. Similar to the motion compression network, the context compression network also consists of an encoder, an entropy model, and a decoder, which together generate the latent feature pair for compression and the intermediate feature pair at different scales during feature extraction and reconstruction process.
Frame Reconstruction. The reconstructed feature pair $\{\bm{F}_{t}^{L},\bm{F}_{t}^{R}\}$ is fed into the frame reconstruction network to generate the reconstructed frame pair $\{\bm{\hat{x}}_{t}^{L},\bm{\hat{x}}_{t}^{R}\}$ at the current time step.
Among the five components described above, the motion estimation and motion compensation modules primarily leverage temporal information to reduce the transmission bitstream, whereas the frame reconstruction module serves as a general component for recovering high-quality images from the processed features. To further address the cross-view redundancy inherent in stereo videos, this work aims to enhance the quality of the quantized latent feature pair and intermediate feature pair through a carefully designed disparity compensation mechanism.
III-B Hybrid Disparity Compensation
As mentioned earlier, our goal is to improve the quality of intermediate and quantized latent features for both the left and right views during the motion and context compression processes. To this end, we exploit cross-view correlations to enhance feature representations. Specifically, in both the motion and context compression modules, we align the intermediate or latent features of the left (resp., right) view with those from the right (resp., left) view.
To enable effective cross-view feature alignment, we employ our Hybrid Disparity Compensation (i.e., HDC) strategy. HDC is integrated into three key components of each module: the multi-scale encoder, the entropy model, and the multi-scale decoder, as illustrated in the central part of Fig. 1111For simplicity, we denote the multi-scale intermediate feature pairs produced by the encoder and decoder as $\{K_{t}^{L},K_{t}^{R}\}$, and the latent feature pairs as $\{y_{t}^{L},y_{t}^{R}\}$, for both the motion and context compression modules.. First, to produce higher-quality intermediate feature pairs $\{K_{t}^{L},K_{t}^{R}\}$, we propose Hybrid Disparity Compensation for Feature Extraction and Reconstruction (i.e., HDC-FER). This module aligns and fuses cross-view features during both the feature extraction and reconstruction stages, thereby enhancing the quality of the resulting features. Second, to improve the latent feature pairs $\{y_{t}^{L},y_{t}^{R}\}$ for more effective entropy modeling in both views, we introduce Hybrid Disparity Compensation for Entropy Modeling (i.e., HDC-EM). This technique aligns cross-view features and leverages them as prior information to enhance the accuracy of entropy estimation.
In general, both HDC-FER and HDC-EM are specialized instantiations of our high-level HDC method, as illustrated in the right part of Fig. 1 Specifically, HDC combines explicit pixel displacement, providing a robust structural prior with implicit cross-attention mechanisms for effective disparity compensation. To capture explicit pixel displacement, we horizontally shift features from both views to construct feature volumes across multiple disparity levels. A similarity map is then computed using an element-wise dot product between these volumes, followed by normalization to produce an attention score. This score is subsequently utilized in a cross-attention mechanism to implicitly align features across views. Detailed descriptions of HDC-FER and HDC-EM are provided in Sec.III-C1 and Sec. III-C2, respectively.
III-C Hybrid Disparity Compensation for NSVC
III-C1 HDC for Feature Extraction and Reconstruction (HDC-FER)
In NSVCs, the encoder and decoder serve crucial roles in transforming the raw input and its latent representation into each other, effectively handling feature extraction and reconstruction. A sophisticated latent representation not only reconstructs the input with high fidelity but also remains amenable to compression, underscoring the importance of advanced encoder–decoder architectures in neural-based compression models. In this subsection, we explain how our Hybrid Disparity Compensation module for Feature Extraction and Reconstruction (HDC-FER) leverages cross-view information in features of both views to reduce cross-view redundancy during encoding and produce more informative features during decoding, thereby enhancing stereo video compression efficiency.
Our HDC-FER module is designed to enhance feature representation during feature extraction and reconstruction by leveraging HDC-based cross-view alignment. Specifically, HDC-FER performs bidirectional alignment, mapping features from each view to its counterpart, enabling mutual enhancement and improving stereo video compression efficiency. For simplicity, we denote the intermediate features of the left view and right view in the encoder-decoder as $\bm{K}^{L}$ [^2] For brevity, the time index $t$ is omitted from all variables in Sec. III-C, including variables shown in figures.. and $\bm{K}^{R}$, respectively (see Fig. 2). Since there exists a significant disparity offset along the width dimension between $\bm{K}^{L}$ and $\bm{K}^{R}$, it is hard for the convolutional networks to aggregate cross-view information for the two views due to a limited receptive field. To address this, we follow prior stereo-processing works [7, 3, 5] and shift feature maps horizontally from 1 to a maximum disparity value, constructing 4D feature volumes $\bm{V}^{L}$ and $\bm{V}^{R}$ with the shape of $D\times C\times H\times W$:
$\begin{split}\bm{V}^{L}(d,c,h,w)&=\bm{K}_{\downarrow}^{L}(c,h,w+d),\\ \bm{V}^{R}(d,c,h,w)&=\bm{K}_{\downarrow}^{R}(c,h,w-d),\end{split}$ (1)
where $C$, $H$ and $W$ denote the number of channels, height, and width of the feature maps, and the corresponding lowercase letters are their respective indices; $\bm{K}^{L}_{\downarrow}$ and $\bm{K}^{R}_{\downarrow}$ are the downsampled intermediate feature maps as shown in Fig. 2. Next, we compute a cross-view similarity map using an element-wise dot product between the feature volumes:
$\begin{split}\bm{F}=\bm{V}^{L}\odot\bm{V}^{R}.\end{split}$ (2)
Figure 2: The proposed Hybrid Disparity Compensation Module for Feature Extraction and Reconstruction (HDC-FER). We input the intermediate feature pair $\{\bm{K}^{L},\bm{K}^{R}\}$ produced by the encoder and decoder modules from Motion and Context Compression (See Fig. 1). We first perform downsampling and horizontally shifting operation for the intermediate feature $\bm{K}^{L}$ (resp., $\bm{K}^{R}$) to obtain the 4D disparity volume $\bm{V}^{L}$ (resp., $\bm{V}^{R}$), where we shift the $\bm{K}^{L}$ (resp., $\bm{K}^{R}$) across a disparity range from 11 to $D$ and $D$ denotes the maximum disparity value. Subsequently, an attention score $\bm{F}^{*}$ (i.e., similarity map) is generated by applying standard operations, element-wise dot product, $\operatorname{Softplus}$ function and $\operatorname{Tanh}$ function. Then, we apply this score into the previously produced cost-volume $\bm{V}^{L}$ (resp., $\bm{V}^{R}$). After a refinement, we can obtain final aligned reference feature $\bm{K}_{Ref}^{R}$ (resp., $\bm{K}_{Ref}^{L}$), which will be added back to the intermediate feature $\bm{K}^{L}$ (resp., $\bm{K}^{R}$) to produce new intermediate feature $\bm{K}_{new}^{L}$ (resp., $\bm{K}_{new}^{L}$).
This similarity map $\bm{F}$ is then normalized using the $\operatorname{Softplus}$ and $\operatorname{Tanh}$ functions, inspired by the $\operatorname{Mish}$ activation function [56], to suppress irrelevant values and enhance significant correlations. The resulting attention score $\bm{F}^{*}$ serves as a disparity-aware cross-view similarity measure, indicating how well features at different disparity shifts correspond across views. To aggregate aligned cross-view features, we employ a cross-attention mechanism using a weighted soft-warp operation. Specifically, we perform a 3D convolution over the weighted feature volumes to obtain reference features $\bm{K}^{L}_{\rm{Ref}}$ and $\bm{K}^{R}_{\rm{Ref}}$:
[…]
These reference features undergo upsampling and are added to the original inputs, enabling each view to enhance its feature representation using complementary information from the other view. Thus, HDC-FER integrates the HDC-based cross-attention mechanism into feature extraction and reconstruction, ensuring that each view benefits from the most relevant cross-view features. During encoding, it helps both views reduce cross-view redundancy, thereby improving compression efficiency. During decoding, it yields a more informative and compact latent representation that further improves the reconstruction quality of stereo video frames.
III-C2 HDC for Cross-View Entropy Modeling (HDC-EM)
In stereo video compression, entropy modeling refers to modeling the probability distribution of latent features for arithmetic coding (the lossless compression stage of a codec). A more accurate entropy model yields a better estimate of the latent feature distribution, thereby reducing the required bitstream size. In this work, we propose a Hybrid Disparity Compensation for Cross-View Entropy Modeling (HDC-EM) module to improve entropy prediction accuracy by exploiting cross-view correlations. HDC-EM enhances context-based entropy coding by partitioning quantized latent features of each view into multiple channel-wise slices for progressive encoding. During this progressive encoding, previously encoded latent features from both views serve as priors. The HDC mechanism aligns the opposite view’s features to the current view, refining these cross-view priors. This cross-view alignment provides more informative priors for the entropy model, enabling more accurate probability predictions and ultimately more efficient compression.
Figure 3: (a). Overview of the Hybrid Disparity Compensation module for Entropy Modeling (HDC-EM). The input is the latent feature pair $\{\bm{y}^{L},\bm{y}^{R}\}$ produced by the encoder modules of the Motion and Context Compression components (see the middle part of Fig. 1). Each latent feature $\bm{y}^{L}$ (resp., $\bm{y}^{R}$) is evenly divided along the channel dimension into N slices $\bm{y}_{1\sim N}^{L}$ (resp., $\bm{y}_{1\sim N}^{R}$). Meanwhile, spatial–temporal context features $\bm{\varPhi}_{1\sim N}^{L}$ (resp., $\bm{\varPhi}_{1\sim N}^{R}$) are generated using the network in (c). To estimate the distribution of each quantized latent slice $\bm{\hat{y}}^{L}$ (resp., $\bm{\hat{y}}^{R}$), we use both the context features and the accumulated losslessly compressed slices from the same and opposite views as priors. These are passed to the entropy estimation module $Est_{n}^{M}$ in (b) for accurate entropy coding. (b). The entropy estimation network $Est_{n}^{M}$ predicts the probability distribution for the n-th slice $\bm{\hat{y}}_{n}^{M}$ at the current view $M$, using the accumulated losslessly compressed slices from the same view $M$, (i.e., $\bm{\hat{y}}_{1\sim n-1}^{M}$) and from the opposite view $\overline{M}$ (i.e., $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$, s.t., q=n1q=n-1 for $M=L$ and $q=n$ for $M=R$), along with the accumulated context features $\bm{\varPhi}_{1\sim n}^{M}$. To align the cross-view slices $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$ with the current view, we apply the HDC alignment strategy, which shares a similar architecture to that shown on the right side of Fig. 1. (c) The spatial–temporal prior fusion network generates the context $\bm{\Phi}^{M}$ by fusing the hyperprior, extracted from $\bm{y}^{M}$ via the hyper encoder-decoder ($h_{a}$ and $h_{s}$), with the multi-scale features $\bm{Ctx}^{M}_{1\sim 3}$ (see the left part of Fig. 1). The resulting context is then divided into N channel-wise slices, denoted as $\bm{\Phi}_{1\sim N}^{M}$, following the same slicing operation used for latent features. The context fusion module adopts a structure similar to the temporal context encoder in DCVC-TCM [50], while $h_{a}$, $h_{s}$}, and the prior fusion module consist of several convolutional layers.
Without loss of generality, let $\bm{y}^{L}$ and $\bm{y}^{R}$ denote the latent representations produced by the encoder for the left and right views, respectively, and let $\bm{\hat{y}}^{L}$ and $\bm{\hat{y}}^{R}$ be their quantized forms (coding symbols). Following recent channel-wise entropy modeling approaches in learned image compression [57, 58, 59], we evenly partition $\bm{y}^{L}$ and $\bm{y}^{R}$ along the channel dimension into $N$ slices (see Fig. 3 (a): $\{\bm{y}^{L}_{1},\bm{y}^{L}_{2},\dots,\bm{y}^{L}_{N}\}$ and $\{\bm{y}^{R}_{1},\bm{y}^{R}_{2},\dots,\bm{y}^{R}_{N}\}$. These slices are then encoded sequentially (slice by slice), yielding corresponding quantized slices $\{\bm{\hat{y}}^{L}_{1},\bm{\hat{y}}^{L}_{2},\dots,\bm{\hat{y}}^{L}_{N}\}$ and $\{\bm{\hat{y}}^{R}_{1},\bm{\hat{y}}^{R}_{2},\dots,\bm{\hat{y}}^{R}_{N}\}$. To leverage inter-view dependencies, the encoding follows an alternating order between the two views:
$\bm{\hat{y}}_{1}^{L}\rightarrow\bm{\hat{y}}_{1}^{R}\rightarrow\bm{\hat{y}}_{2}^{L}\rightarrow\bm{\hat{y}}_{2}^{R}\rightarrow\cdots\rightarrow\bm{\hat{y}}_{N}^{L}\rightarrow\bm{\hat{y}}_{N}^{R}~.$ (4)
In this progressive coding scheme, each new slice’s entropy modeling is conditioned on previously encoded information. Specifically, for the input slice $\bm{y}_{n}^{M}$ (where $M\in\{L,R\}$, denotes the current view and $\overline{M}$ denotes the opposite view), the entropy estimation network $Est_{n}^{M}$ integrates three types of priors to predict the probability distribution of the quantized slice $\bm{\hat{y}}_{n}^{M}$. First, intra-view priors, which include previously encoded slices from the same view, $\bm{\hat{y}}_{1\sim n-1}^{M}=\{{\bm{\hat{y}}_{1},\dots,\bm{\hat{y}}_{n-1}}\}^{M}$. These capture the already-known information from earlier slices of the current view. Second, cross-view priors, which consist of encoded slices from the opposite view, $\bm{\hat{y}}_{1\sim q}^{\overline{M}}=\{\bm{\hat{y}}_{1},\dots,\bm{\hat{y}}_{q}\}^{\overline{M}}$, where q=n1q=n-1 for $M=L$ (i.e., the left view) and $q=n$ for $M=R$ (i.e., the right view). In other words, when encoding a left-view slice, we use all slices up to index n1n-1 from the right view as cross-view priors, and for a right-view slice we use up to index $n$ from the left view. The HDC mechanism is applied to these cross-view priors to align them with the current view, compensating for disparity via disparity-aware shifting and cross-attention fusion. Third, spatial-temporal priors, which capture spatial and temporal context for the current slice, are denoted as $\bm{\Phi}_{1\sim n}^{M}$. These priors are obtained by fusing the hyperprior features with multi-scale context features $\bm{Ctx}_{1\sim 3}^{M}$ [^3] For the motion compression network, only hyperprior features are used; context features are not involved., which are derived from a motion compensation module (see the left part of Fig. 1), as illustrated in Fig. 3 (c). The spatial–temporal priors provide additional cues from the broader image context and temporal reference frames. By incorporating these three sets of priors, $Est_{n}^{M}$ can more accurately estimate the probability distribution for $\bm{\hat{y}}_{n}^{M}$, leading to improved entropy modeling for that slice.
The entropy estimation process within $Est_{n}^{M}$ is illustrated in Fig. 3 (b). First, the HDC mechanism performs cross-view alignment: it takes the available encoded slices from the opposite view $\bm{\hat{y}}_{1\sim q}^{\overline{M}}$ and aligns them to the current view $M$ using disparity compensation. This alignment uses spatial–temporal prior of the current view as an anchor to guide the disparity-aware shifting and cross-attention, similar to Sec. III-C1. In contrast to the previous HDC-FER module, this alignment step simplifies the process by excluding any downsampling, upsampling, or additive operations, and directly yields the aligned cross-view priors (denoted as $\bm{s}_{1\sim q}^{\overline{M}}=\{\bm{s}_{1},\dots,\bm{s}_{q}\}^{\overline{M}}$).
Figure 4: Rate-distortion (RD) curves. The results are evaluated on the KITTI 2012, KITTI 2015 and Nagoya [21] datasets in terms of Bpp-PSNR.
Next, the entropy estimation network fuses the aligned cross-view priors with the intra-view and spatial–temporal priors for the current slice. In particular, $\bm{s}_{1\sim q}^{\overline{M}}$, $\bm{\hat{y}}_{1\sim n-1}^{M}$, and $\bm{\Phi}_{n}^{M}$ are concatenated and processed by a dedicated convolutional module (a DepthConvBlock). This module outputs the parameters of a parametric probability distribution for the current latent slice, mean $\bm{\mu}_{n}^{M}$ and scale $\bm{\sigma}_{n}^{M}$. Using these predicted parameters, the encoder then quantizes the input slice $\bm{y}_{n}^{M}$ by subtracting the mean, rounding to the nearest integer, and adding the mean back:
$\bm{\hat{y}}_{n}^{M}=\Big\lfloor\bm{y}_{n}^{M}-\bm{\mu}_{n}^{M}\Big\rceil+\bm{\mu}_{n}^{M}~,$ (5)
where $\lfloor\cdot\rceil$ denotes rounding to the nearest integer. The resulting quantized slice $\bm{\hat{y}}_{n}^{M}$ is stored and treated as a known prior for subsequent entropy estimation steps, enabling the encoding process to adapt progressively as more slices (and cross-view information) become available.
Overall, by integrating HDC-based cross-view alignment into the entropy model, the HDC-EM module yields more accurate probability estimates for each latent slice. This improvement in entropy modeling translates to fewer bits required for encoding, thereby enhancing overall compression efficiency.
III-D Objective Function
We set the rate-distortion cost sum of the two views as optimization target like formal methods:
$\mathcal{L}_{t}=\sum_{M\in\{L,R\}}\{\lambda d(\bm{x}_{t}^{M},\bm{\hat{x}}_{t}^{M})+r(\bm{\hat{y}}_{t}^{M})+r(\bm{\hat{z}}_{t}^{M})\},$ (6)
where $M$ denotes the view, $L$ or $R$, $d(x_{t}^{M},x_{t}^{M})$ denotes the distortion between the original uncompressed frame $\bm{x}_{t}^{M}$ and the reconstruction frame $\bm{\hat{x}}_{t}^{M}$ at time step $t$, which is usually measured by MSE or MS-SSIM loss. The term $r(\bm{\hat{y}}_{t}^{M})+r(\bm{\hat{z}}_{t}^{M})$ represents the total bit rate for representing transmission features in both motion and context compression. $\lambda$ is a hyperparameter used to control the trade-off factor between the total bit rate and distortion, controlling the balance between transmitted information and reconstruction quality.
IV Experiments
IV-A Evaluation Datasets.
For evaluation, previous NSVC approaches [4, 38] primarily focused on autonomous driving datasets, such as Cityscapes [60] and KITTI [19, 20], limiting their assessments to this specific domain. To ensure a fair comparison, we follow their experimental settings for Cityscapes and KITTI, adopting the same frame selection and preprocessing. Furthermore, to evaluate performance beyond autonomous driving scenarios, we extend our experiments to the Nagoya dataset [21], which represents more general multi-view scenes.
[…]
Figure 5: Subjective quality comparison on KITTI 2012 dataset.
[…]
Nagoya. In contrast to Cityscapes and KITTI, which focus on autonomous driving scenarios, the Nagoya dataset [21] provides multi-view video sequences of general scenes. To assess the generalization capability of our model, we selected two standard MV-HEVC benchmark sequences, Kendo and Balloons, available through the Nagoya MPEG-FTV project [21]. Evaluation was performed on the first 96 frames, utilizing the first and third views from both sequences. The frames were kept at their original 1024 × 768 resolution for testing. The GOP was set to 32 for evaluation. We test on the RGB videos which are obtained from the original YUV420 format with BT.709 conversion.
IV-B Experiment Protocols
Figure 6: Ablation study on rate-distortion (RD) performance by progressively integrating the proposed HDC-EM and HDC-FER modules into the baseline DCVC-TCM*.
Evaluation Metrics. To assess the quality of frame reconstruction, we adopt RGB-PSNR as the distortion metric. Compression efficiency is evaluated using bits per pixel (Bpp). Furthermore, BD-rate [61] is employed to provide a comprehensive quantitative comparison. For the initial frames of both views (i.e., the I-frames), compression is performed independently using the learned image compression model proposed by He et al. [58], while the proposed method is applied to the remaining frames (i.e., the P-frames).
TABLE I: BD-RATE (%) COMPARISON IN RGB COLORSPACE MEASURED WITH PSNR. THE ANCHOR IS MV-HEVC.
Method Cityscapes KITTI 2012 KITTI 2015 Kendo Balloons
MV-HEVC [2] 0.0 0.0 0.0 0.0 0.0
HEVC [23] 34.14 7.77 12.64 26.28 26.02
FVC [27] -14.9 -2.35 0.97 76.66 126.84
DCVC-TCM* -24.17 -48.61 -46.82 37.02 66.67
LSVC [4] -32.05 -17.18 -13.47
LLSS [3] -49.44 -18.18 -15.76
Ours -53.13 -55.97 -54.72 -19.52 -10.61
Implementation Details. We train all models by minimizing a rate-distortion objective, weighted by a Lagrange multiplier $\lambda$, as formulated in Eq. 6. Following [3], we set the maximum disparity value $D$ to 192, corresponding to the maximum disparity reported in the KITTI datasets [19, 20]. All models are implemented in PyTorch and optimized with the Adam optimizer[62].
To ensure stable optimization, we divide the training process into four stages. In Stage 1, following the training process used in DCVC-TCM [50], a single-view variant of our model (i.e., DCVC-TCM*)—excluding all HDC-FER modules and cross-view priors in HDC-EM entropy model—is pretrained on the Vimeo-90K dataset [63] with a batch size of 4. Subsequently, the model is fine-tuned on the Cityscapes dataset [60] with a fixed learning rate of $1\times 10^{-5}$ during Stages 2 to 4.
Figure 7: Ablation study on RD performance by removing the attention score and shift operation from the proposed HDC mechanism.
In Stage 2, cross-view priors are introduced into HDC-EM entropy model, and the network is fine-tuned for 300k iterations. In Stage 3, HDC-FER modules are added and trained for 20k iterations, with all other parameters frozen. Finally, in Stage 4, the complete model (including both HDC-EM and HDC-FER) undergoes an additional 200k iterations of fine-tuning.
All training is conducted on two NVIDIA RTX 4090 GPUs. During training, input frames are cropped to $256\times 256$ in Stage 1 and $256\times 384$ in the subsequent stages.
IV-C Comparison with State-of-the-art Methods
Baseline Methods. Following [3, 4], we also compare our method against traditional video codecs HEVC [23], a traditional MVC method MV-HEVC [2], a neural video codec FVC [27] and two state-of-the-art NSVC methods LSVC [4] and LLSS [3]. Specifically, we adopt HM-16.20 [64] and HTM-16.3 [65] as the official reference implementations of HEVC and MV-HEVC following [4, 3], respectively. For HEVC, we use the encoder_lowdelay_P_main configuration, and for MV-HEVC, we adopt the baseCfg_2view setting. The results of LSVC [4] and LLSS [3] are reported from their original papers.
In addition, we compare our approach with DCVC-TCM*, a lightweight variant of DCVC-TCM [50], which serves as the backbone of our framework. DCVC-TCM* reduces model complexity by employing fewer channels and a lightweight motion estimation module [66], while maintaining the overall architecture. These modifications significantly lower the computational cost and reduce encoding latency, making it more suitable for resource-constrained environments. As shown in Tab. II, DCVC-TCM* achieves substantial improvements in computational efficiency and encoding speed. Our framework extends DCVC-TCM* by incorporating HDC-FER and HDC-EM modules to further enhance compression efficiency. For fair comparison, DCVC-TCM* is trained following the similar pipeline as our method, with pretraining on Vimeo-90K [63] and fine-tuning on Cityscapes [60].
TABLE II: Model complexity on $512\times 512$ stereo videos, evaluated under the settings of [3]. DCVC-TCM* refers to a lightweight version of DCVC-TCM [50], used as our baseline.
Method Params (M) FLOPs (G) Enc Speed Dec Speed
DCVC-TCM [50] 10.55 735.95 14.29fps 17.85fps
DCVC-TCM* 9.68 431.24 18.87fps 18.52fps
LSVC [4] 50.50 1760.0 6.25fps 8.33fps
LLSS [3] 39.22 634.9
Ours 16.39 536.90 8.33fps 10.00fps
Rate-Distortion Performance. The rate–distortion (RD) curves in Fig. 4 compare our proposed method with baseline approaches in terms of Bpp–PSNR across four datasets: KITTI 2012 [19], KITTI 2015 [20], and two general-scene video sequences, Kendo and Balloons, from the Nagoya dataset [21]. As shown in the figure, our method consistently achieves state-of-the-art performance across all datasets, demonstrating substantial improvements in RD efficiency over both traditional and neural-based baselines.
Table I presents a detailed quantitative analysis using BD-Rate [61] as the evaluation metric, computed based on Bpp–PSNR with MV-HEVC [2] serving as the anchor. On autonomous driving datasets (i.e., Cityscapes, KITTI 2012, and KITTI 2015), our method achieves an average BD-rate reduction of 54.61% compared to MV-HEVC and 26.81% relative to the previous state-of-the-art NSVC method LLSS [3].
Moreover, on general-scene multi-view video sequences (i.e., Kendo and Balloons), our approach attains an average BD-rate saving of 15.07% compared to MV-HEVC, and surpasses the single-view baseline DCVC-TCM* by 66.91% on average. These results highlight our model’s superior capability in leveraging disparity redundancy, particularly in indoor stereo scenes—an area where prior NSVC methods such as LSVC and LLSS have primarily demonstrated benefits only in driving scenarios.
Overall, the experimental results consistently demonstrate the superiority of our method over both traditional codecs and recent learning-based stereo video compression frameworks, achieving lower BD-rate values across all evaluated datasets.
Subjective quality comparison. Fig. 5 presents the subjective quality comparison. Compared with MV-HEVC [2], the proposed method achieves noticeably better visual quality at comparable or lower bit rates, and preserves more fine details.
IV-D Computational Complexity
Table II compares the FLOPs, parameter counts (i.e., Params), and latency (i.e., Enc Speed and Dec Speed) of our model with DCVC-TCM [50], DCVC-TCM*, LSVC [4], and LLSS [3], on a $\times512$ input frame pair, following the evaluation protocol described in [3]. The FLOPs and parameter counts for LLSS are cited from its published results. We measure FLOPs, parameters, and latency for DCVC-TCM, DCVC-TCM*, LSVC, and our model using an NVIDIA RTX 4090 GPU.
As shown in Table II, DCVC-TCM*, a simplified variant of the backbone model DCVC-TCM, reduces the FLOPs from 735G to 431G, significantly lowering computational complexity. Incorporating the proposed HDC-FER and HDC-EM modules into DCVC-TCM* moderately increases the FLOPs to 536.90G.
Compared to previous state-of-the-art NSVC methods such as LSVC and LLSS, our approach achieves better compression performance while maintaining significantly lower computational overhead, fewer parameters, and faster encoding and decoding speeds (compared to LSVC). These results validate the efficiency and effectiveness of the proposed modules, making our method highly suitable for deployment in resource-constrained environments.
IV-E Ablation Study
Effectiveness of HDC-FER and HDC-EM. To evaluate the individual and combined effectiveness of our proposed HDC-FER and HDC-EM modules, we conduct an ablation study based on the single-view baseline model DCVC-TCM* using the Cityscapes dataset. Specifically, the variant DCVC-TCM* + HDC-FER integrates HDC-FER modules into the feature extraction and reconstruction process of the baseline method DCVC-TCM*, enabling enhanced feature representation during the encoding and decoding stages to simultaneously compress the two views of stereo frames. DCVC-TCM* + HDC-EM replaces the original single view entropy module of DCVC-TCM* with the proposed cross-view HDC-EM entropy model, which leverages information from both views to achieve more accurate entropy estimation for each view in an autoregressive manner.
The results presented in Fig. 6 demonstrate that integrating HDC-FER alone (i.e., DCVC-TCM* + HDC-FER) achieves a BD-rate reduction of 6.03%, and substituting the entropy module with HDC-EM (i.e., DCVC-TCM* + HDC-EM) results in a BD-rate reduction of 26.59%. When both modules are simultaneously integrated into the model (i.e., DCVC-TCM* + HDC-EM + HDC-FER), the total BD-rate reduction is further improved to 32.17%, thereby confirming the significant benefits of employing cross-view information and enhanced feature representation for compression performance.
Figure 8: Ablation study on rate-distortion (RD) performance with various single-view NVC backbone networks in our framework.
Effectiveness on Implicit and Explicit Compensation in HDC. To validate the effectiveness of the implicit and explicit compensation within our HDC mechanism, we conduct an ablation study on the Cityscapes dataset, focusing on the “attention score” (i.e., implicit cross-attention) and the shift operation (i.e., explicit pixel displacement). First, removing the “attention score” from both HDC-EM and HDC-FER modules results in cross-view features being extracted solely from the shifted feature volumes, without the guidance of attention weighting. As shown in Fig. 7, this modification leads to a 13.75% increase in BD-rate, highlighting the significant role of the implicit cross-attention mechanism in improving compression efficiency. Next, to assess the contribution of explicit pixel displacement, we disable the shift operation in both modules, thereby allowing feature interaction between views only through simple channel attention. This results in a 9.08% increase in BD-rate, as illustrated in Fig. 7, demonstrating the effectiveness of explicit pixel displacement in facilitating cross-view feature extraction.
Effectiveness of Single-View NVC Backbone. To rigorously assess the effectiveness of the proposed HDC-FER and HDC-EM modules, we perform a controlled comparison on the KITTI 2012 dataset by integrating them into the same backbone as prior SOTA NSVC methods, including LSVC [4] and LLSS [3]. We also attempted to incorporate our modules with other state-of-the-art NVC backbones, such as DCVC-DC [41] and DCVC-FM [43]; however, the pretrained weights provided by these methods are highly sensitive to fine-tuning, and even using a very small learning rate results in a severe drop in performance. Therefore, for a comprehensive evaluation, we adopt the official pretrained FVC [27] model trained on the Vimeo-90K dataset [63], integrate our modules into its framework, and finetune the resulting model on Cityscapes [60], following the protocol detailed in Sec. IV-B.
As presented in Fig. 8, our method achieves superior performance over prior SOTA approaches under identical backbone settings, thereby substantiating the contribution of the proposed modules to compression efficiency. Moreover, when employing DCVC-TCM* as the backbone—which corresponds to the final configuration of our model—further performance gains are observed, reaffirming the general applicability of the proposed HDC-FER and HDC-EM modules.
TABLE III: Performance comparison of cross-view interaction strategies on BD-rate (%) and latency. The proposed method is used as the anchor for BD-rate calculation.
Interaction Strategy BD-Rate (%) Latency (ms)
Stereo Attention [13] 8.19 228.54
Mutual Attention [67] 7.03 224.45
HDC (proposed) 0.0 158.26
$\ddagger$} Latency denotes the average inference time of a forward pass, measured on a pair of $512\times 512$ stereo frames, following the setting in [3].
Cross-view Interaction Strategy Comparison. To verify the effectiveness of the proposed Hybrid Disparity Compensation (HDC) mechanism, we conduct a comprehensive comparison against state-of-the-art cross-view interaction strategies in the context of stereo image compression. Specifically, we replace the cross-view interaction component within both the HDC-EM and HDC-FER modules with the Stereo Attention mechanism [13] and the Mutual Attention mechanism [67], respectively, while keeping all other components unchanged to ensure a fair comparison. This ablation study jointly evaluates compression performance and inference latency, thereby providing a holistic assessment of the proposed HDC mechanism.
As presented in Table III, replacing the HDC mechanism with Stereo Attention results in a performance degradation of 8.19%, while substituting it with Mutual Attention leads to a 7.03% drop. Moreover, our HDC mechanism also achieves significantly lower inference latency, reducing it by approximately 30% compared to both Stereo Attention and Mutual Attention mechanisms.
V Conclusion and Future Work
In this work, we proposed a neural stereo video compression framework that incorporates a novel hybrid disparity compensation strategy to effectively reduce cross-view redundancy. By integrating explicit structural priors with implicit disparity alignment, our method achieves a synergistic balance between efficiency and robustness. Extensive experiments demonstrate that our approach achieves state-of-the-art compression performance on both automotive driving and standard indoor stereo benchmarks, outperforming existing traditional and neural stereo video compression methods. Ablation studies further validate the effectiveness and efficiency of our proposed modules. Beyond stereo video, the proposed technique shows strong potential for multi-view applications, where specifically designed cross-view interaction orders enable effective exploitation of inter-view redundancies, representing a promising direction for future research.